"During a recent year in Alaska<5238 boys and 4984 girls were born. if a new born is selected from that year, what is the probability that the baby is a girl."
"

.

Answer by flame8855(424)

total = 5238 + 4984
10222
probability = 4984/10222 = 0.4875